The length of a rectangle is 1 less than three times it's width. The perimeter is 30 feet. Find the length of this rectangle
Answer by stanbon(75887)

The length of a rectangle is 1 less than three times it's width. The perimeter is 30 feet. Find the length of this rectangle
====
Let width be "x".
Then length = "3x-1".
----
Equation:
Perimeter = 2(length + width)
---
30 = 2(3x-1+x)
---
15 = 4x-1
4x = 16
x = 4
----
length = 3x-1 = 3*4-1 = 11 feet
====================================
Cheers,
Stan H.
============